▼ is based on the notion that a dollar paid in the future is less valuable than a dollar paid today The present value of a loan in which $1000 is to be paid out a year from today with the interest rate equal to 1% is S decimal place) Round your response to the neareast two If a loan is paid after two years, and the amount S700 response to the neareast two decimal place) is to be paid hen with a corresponding 2% interest rate, the present value of the loan is S Round your
Answer #1

Sol:- Present value

990 (It is only for a year so simple interest will be deducted)

686.207

(pv=700/(1+0.01)2 =686.207)

PV=FV/(1+r) n

r=rate of interest, n=no. Of year